OpenStudy (anonymous):

If an object is dropped from a height of 144 feet,the function h(t)= - 16t^2 + 144 gives the height of the object after t seconds. When will the object hit the ground?

OpenStudy (anonymous):

AT GROUND H=O =-16T^2+144 T=SQRT(144/16)T

OpenStudy (anonymous):

Since it hits the ground, h(t) =0 ( that is it is o the ground..) h(t)= 0 - 16t^2 + 144 =0 Subtract both sides by 144, you'll get -16t^2 = -144 Then, divide both sides by -16, you'll get t^2 = 9 Now, take square root for both sides, what do you get?

OpenStudy (anonymous):

its 3 s

OpenStudy (anonymous):

Yes!~